# LUCID: LLM-Generated Utterances for Complex and Interesting Dialogues

### v1

@@ Appendix F Slot duplication within PRESTO and SGD
Appendix G Example Dialogues
In addition to the examples provided in Figure 1, Figure 3 and Figure 4, we provide three additional examples of the LUCID generated conversations. To provide an unbiased sample of our conversations, we show the first three dialogues in the dataset (see Figure 5, Figure 6 and Figure 7).
Appendix H Breakdown of issues identified by the LLM validation stages
In total, the two LLM validation agents described in stage 13 identified issues for 2,457 and 2,477 of turns, compared to the 557 issues identified by the LLM validation agent in stage 14 with access to the conversation rules. While the agent with access to these rules identifies fewer errors, a larger proportion of the issues identified are not identified by another validation agent (23%, compared to 7% for each other validation agent).
Appendix I Modelling setup, parameters, computing setup
For each baseline experiment, we train for 3 epochs. This was selected as a hyper-parameter based on the development set loss after training for 1-10 epochs for our Flan-T5-base baseline. We use a learning rate of $\times 10^{-5}$, with a linear learning schedule, a batch size of 2 with 8 gradient accumulation steps. For experiments with reduced training data, we train for more epochs as the training data is increased (with epochs inversely proportional to the size of the training data, allowing for a fair comparison).
@@ Appendix I Modelling setup, parameters, computing setup
Figure 6: An example conversation from LUCID (Example #2). As described in Appendix G, we show the first three examples to provide an unbiased sample from our generated data.
Figure 7: An example conversation from LUCID (Example #3). As described in Appendix G, we show the first three examples to provide an unbiased sample from our generated data.
@@
Table 4: Reported statistics for LUCID, and related datasets for task-oriented dialogue. All statistics refer to the training split of the datasets, except for Frames which reports figures for all splits. Compared to previous dialogue datasets, LUCID has considerably more slots, and more turns per dialogue. There are also more possible slot values per turn than either MultiWoZ or SGD.
@@
Table 5: Results of our baseline model trained for 3 epochs, using a SentenceBERT retrieval model. Each evaluation metric is described in more detail in Appendix C. Results are from a single seed in each case.
@@
Table 6: Results of a T5-Flan-base model with our tool retrieval, using oracle tools, and with no tools provided in the prompt. Each evaluation metric is described in more detail in Appendix C. Results are from a single seed in each case.
@@
Table 7: Results of a T5-Flan-base model trained with varying amounts of training data (count of the system turns provided). Each evaluation metric is described in more detail in Appendix C.

### v2

@@ Appendix I Slot Duplication within PRESTO and SGD
Appendix J Example Dialogues
In addition to the examples provided in Figure 1 and Figure 3, we provide three additional examples of the LUCID generated conversations. To provide an unbiased sample of our conversations, we show the first three dialogues in the dataset (see Figure 5, Figure 6 and Figure 7). We also show examples of each of the unhappy paths used in our dataset (see Figure 4).
Appendix K Modelling Setup, Parameters, Computing Setup
For each baseline experiment, we train for 3 epochs. This was selected as a hyper-parameter based on the development set loss after training for 1-10 epochs for our Flan-T5-base baseline. We use a learning rate of $\times 10^{-5}$, with a linear learning schedule, a batch size of 2 with 8 gradient accumulation steps. For experiments with reduced training data, we train for more epochs as the training data is increased (with epochs inversely proportional to the size of the training data, allowing for a fair comparison). For each LLM call, we use a temperature value of 0.7.
@@
Figure 6: An example conversation from LUCID (Example #2). As described in Appendix J, we show the first three examples to provide an unbiased sample from our generated data.
Figure 7: An example conversation from LUCID (Example #3). As described in Appendix J, we show the first three examples to provide an unbiased sample from our generated data.
@@
Table 5: Reported statistics for LUCID, and related datasets for task-oriented dialogue. All statistics refer to the training split of the datasets, except for Frames which reports figures for all splits. Compared to previous dialogue datasets, LUCID has considerably more slots, and more turns per dialogue. There are also more possible slot values per turn than either MultiWoZ or SGD. The number of turns in LUCID refer to User, System, Signal and Response turns.
@@
Table 6: Results of our baseline model trained for 3 epochs, using a SentenceBERT retrieval model. Each evaluation metric is described in more detail in Appendix G. Results are from a single seed in each case.
@@
Table 7: Results of a T5-Flan-base model with our tool retrieval, using oracle tools, and with no tools provided in the prompt. Each evaluation metric is described in more detail in Appendix G. Results are from a single seed in each case.
@@
Table 8: Results of a T5-Flan-base model trained with varying amounts of training data (count of the system turns provided). Each evaluation metric is described in more detail in Appendix G.